OpenStudy (anonymous):

The area of a rectangle is equal to the product of the length of the rectangle times its width. Find the area of a rectangular field that has a length, l, of 2 4/5 km and a width, w, of 3 1/8 km. Express your answer as a mixed number or improper fraction in lowest terms. ? km2

ganeshie8 (ganeshie8):

multiply both the numbers to get the area

ganeshie8 (ganeshie8):

Area = length x width = (2 4/5) * (3 1/8)

ganeshie8 (ganeshie8):

2 4/5 is same as 2+4/5

ganeshie8 (ganeshie8):

3 1/8 is same as 3+1/8

ganeshie8 (ganeshie8):

\[\large \left(2+\frac{4}{5}\right) * \left(3+\frac{1}{8}\right)\]

OpenStudy (anonymous):

5 4/40

ganeshie8 (ganeshie8):

get common denominator and add the fractions first

ganeshie8 (ganeshie8):

\[\large \left(2+\frac{4}{5}\right) * \left(3+\frac{1}{8}\right)\] \[\large \left(\frac{2*5}{5}+\frac{4}{5}\right) * \left(\frac{3*8}{8}+\frac{1}{8}\right)\] \[\large \left(\frac{10}{5}+\frac{4}{5}\right) * \left(\frac{24}{8}+\frac{1}{8}\right)\] \[\large \left(\frac{14}{5}\right) * \left(\frac{25}{8}\right)\]

OpenStudy (anonymous):

350/40

ganeshie8 (ganeshie8):

yes 35/4

ganeshie8 (ganeshie8):

change it to mixed number

OpenStudy (anonymous):

8 3/4

ganeshie8 (ganeshie8):

Correct !

OpenStudy (anonymous):

thank you soo much

ganeshie8 (ganeshie8):

yw :)